OpenStudy (anonymous):

Is f(x) = 0.69(1.03)x a decreasing or increasing function?

OpenStudy (rea201):

All you need to do for this problem is graph the equation

OpenStudy (anonymous):

ohh thanks

OpenStudy (rea201):

Desmos is a really good online grapher. If you graphed the equation you would see that the equation is increasing.

OpenStudy (rea201):

Also can you give me a medal?

OpenStudy (rea201):

all you have to do is click the best response button

OpenStudy (anonymous):

there you go :)

OpenStudy (rea201):

Thanks!!

OpenStudy (solomonzelman):

no need for graphing

OpenStudy (anonymous):

oh wait by what percent?

OpenStudy (solomonzelman):

\(\large\color{black}{ f(x) = 0.69(1.03)^x}\) when \(\large\color{black}{x=1}\), then you have: \(\large\color{black}{ 0.69 \times1.03}\) when \(\large\color{black}{x=2}\), then you have: \(\large\color{black}{ 0.69 \times 1.03 \times 1.03}\) when \(\large\color{black}{x=3}\), then you have: \(\large\color{black}{ 0.69 \times 1.03 \times 1.03 \times 1.03}\) when \(\large\color{black}{x=4}\), then you have: \(\large\color{black}{ 0.69 \times 1.03 \times 1.03 \times 1.03\times 1.03}\) and on

OpenStudy (solomonzelman):

right?

OpenStudy (anonymous):

yeah

OpenStudy (solomonzelman):

so each time (as \(\large\color{black}{ x}\) increases by \(\large\color{black}{ 1}\) ) you multiply times \(\large\color{black}{ 1.03}\)

OpenStudy (solomonzelman):

that means that the function is increasing.

OpenStudy (solomonzelman):

because anytime, when you multiply a(ny) number \(\large\color{black}{ C}\) by a number greater than 1, then the result will be greater than this number \(\large\color{black}{ C}\) .

OpenStudy (solomonzelman):

So for any exponential function \(\large\color{black}{ y=a(b)^x}\) (when \(\large\color{black}{ a}\) is positive) ~ when \(\large\color{black}{b>0}\) , the \(\large\color{black}{ f(x)}\) is increasing. ~ when \(\large\color{black}{b<0}\) , the \(\large\color{black}{ f(x)}\) is decreasing.

OpenStudy (anonymous):

okay and by what percent is it increasing?

OpenStudy (solomonzelman):

okay, you are multiplying times \(\large\color{black}{ 1.03}\) , which is same as taking \(\large\color{black}{ 103 \text{%}}\) (of the previous input) each time.

OpenStudy (solomonzelman):

\(\large\color{black}{ 100 \text{%}}\) is what you had, and \(\large\color{black}{ +3 \text{%}}\) is the extra part.

OpenStudy (anonymous):

ohh ok ok I get that was helpful thank you!

OpenStudy (solomonzelman):

yw